# Ideal transition systems

### v2

@@ 5 Skew-symmetric matrix Schubert varieties
$\mathsf{MSC}^{\mathrm{ss}}_{w}$ =$=\left\{M\in\mathsf{Mat}^{\mathrm{ss}}_{n\times n}:\text{$\operatorname{rank}M_{[i][j]}=\operatorname{rank}w_{[i][j]}$ for all $(i,j)\in[n]\times[n]$}\right\},$ (5.1)
$\mathsf{MSV}^{\mathrm{ss}}_{w}$ =$=\left\{M\in\mathsf{Mat}^{\mathrm{ss}}_{n\times n}:\text{$\operatorname{rank}M_{[i][j]}\leq\operatorname{rank}w_{[i][j]}$ for all $(i,j)\in[n]\times[n]$}\right\},$
@@ 5.1 Another reverse lexicographic transition system
Unlike the story for ordinary matrix Schubert varieties, both of these containments can be strict [29, Ex. 3.12]. Therefore, to get a skew-symmetric version of Theorem 4.3, we cannot use the right hand side of (5.4) as a skew-symmetric analogue of $J_{w}$.
Instead, we consider a more elaborate construction. Suppose $A=\{a_{0}<a_{1}<\dots<a_{r}\}$ and $B=\{b_{0}<b_{1}<\dots<b_{r}\}$ are two sets of $r+1$ positive integers. Define ;
@@
Then let $A\boxplus B:=\{(i,j)\in A\odot B:i\geq j\}\cup\{(i,j)\in B\odot A:i\geq j\}$ and define
@@
This can be shown in a self-contained way, but the elementary argument is fairly tedious. Alternatively, [29, Lem 3.25] asserts (in a special case) that $u^{\mathrm{ss}}_{AB}$ divides $u_{ij}$ if and only if $\operatorname{adiag}(\mathcal{X}_{AB})$ divides $x_{ij}x_{ji}$, which occurs either when $A=B=\{i,j\}$ or $\{A,B\}=\{\{i\},\{j\}\}$. ∎
@@ 5.2 Finite indexing set
For each $w\in I_{{\textsf{FPF}}}$ define $\mathrm{ss}_{n}(w)\in\mathsf{Mat}^{\mathrm{ss}}_{n\times n}$ to be the matrix whose entry in position $(i,j)$ is $1$ if $i<j=w(i)$, $-1$ if $i>j=w(i)$, or else zero. Then each $\mathsf{B}_{n}$-orbit in $\mathsf{Mat}^{\mathrm{ss}}_{n\times n}$ contains an element of the form $\mathrm{ss}_{n}(w)$ for a unique $w\in I_{{\textsf{FPF}}}^{(n)}$ by results in [4]; it is not hard to show that this holds even when $\mathbb{K}$ is not algebraically closed. Given $w\in I_{{\textsf{FPF}}}^{(n)}$ we are therefore motivated to define
@@
$\mathcal{T}^{\mathrm{ss}}=\left\{(I(\mathsf{MSV}^{\mathrm{ss}}_{w}),J^{\mathrm{ss}}_{w}):w\in\mathsf{Mat}^{\mathrm{ss}}_{n\times n}\right\}=\left\{(I(\mathsf{MSV}^{\mathrm{ss}}_{w}),J^{\mathrm{ss}}_{w}):w\in I_{{\textsf{FPF}}}^{(n)}\right\}$
and the second set is uniquely indexed, since if $v,w\in I_{{\textsf{FPF}}}^{(n)}$ then $I(\mathsf{MSV}^{\mathrm{ss}}_{v})\subseteq I(\mathsf{MSV}^{\mathrm{ss}}_{w})$ if and only if $v\leq w$ in a certain Bruhat order on $I_{{\textsf{FPF}}}$ [29, Prop. 3.11].
@@
As in the classical case, the ideal $I(\mathsf{MSV}^{\mathrm{ss}}_{w})$ is maximal if and only if $\mathsf{MSV}^{\mathrm{ss}}_{w}$ is a point, which occurs only when $\mathsf{MSV}^{\mathrm{ss}}_{w}=\{0\}$. The index $w\in I_{{\textsf{FPF}}}^{(n)}$ for this case is
$0^{\mathrm{ss}}_{n\times n}:=(1,n+1)(2,n+2)\cdots(n,2n)(2n+1,2n+2)(2n+3,2n+4)\cdots\in I_{{\textsf{FPF}}}^{(n)}$
@@
Suppose $w\in I_{{\textsf{FPF}}}^{(n)}$ and $(i,j)\in[n]\times[n]$ has $i>j$. Then $(i,j)\in\mathsf{dom}(w)$ if and only if $u_{ij}\in I(\mathsf{MSV}^{\mathrm{ss}}_{w})$ if and only if $u_{ij}\in J^{\mathrm{ss}}_{w}$.
@@
Suppose $0^{\mathrm{ss}}_{n\times n}\neq w\in I_{{\textsf{FPF}}}^{(n)}$ so that $I(\mathsf{MSV}^{\mathrm{ss}}_{w})$ is not a maximal ideal and $\mathsf{MSV}^{\mathrm{ss}}_{w}\neq\{0\}$. Then $\mathsf{dom}(w)$, which is invariant under transpose, cannot contain every position in $[n]\times[n]$, so there must exist an outer corner $(p,q)$ of $\mathsf{dom}(w)$ with $n\geq p\geq q\geq 1$. As explained in [29, §4.1], all outer corners $(p,q)$ of $\mathsf{dom}(w)$ have $w(p)=q$, so as $w$ has no fixed points we can assume that $p>q$.
@@
It is known that $\varnothing\subsetneq\mathcal{C}^{\mathrm{ss}}(w)\subseteq I_{{\textsf{FPF}}}^{(n)}$ by [29, Lem. 4.8]. Finally, let
@@
The claim that $J^{\mathrm{ss}}_{w}\subseteq\operatorname{init}(I(\mathsf{MSV}^{\mathrm{ss}}_{w}))$ holds by [29, Lem 3.37 and Thm. 3.17]. The other parts of axioms (T1) and (T2) in Definition 3.1 were checked in Sections 5.2 and 5.3.
Fix $w\in I_{{\textsf{FPF}}}^{(n)}$ with $I(\mathsf{MSV}^{\mathrm{ss}}_{w})$ not maximal, and then choose an outer corner $(p,q)$ of $\mathsf{dom}(w)$ with $n\geq p>q\geq 1$. The identities required for axiom (T3) when $(I,J)=(I(\mathsf{MSV}^{\mathrm{ss}}_{w}),J^{\mathrm{ss}}_{w})$ are supplied in [29, Lem. 4.2] if we define $f=u_{pq}\notin I(\mathsf{MSV}^{\mathrm{ss}}_{w})$ and $\Phi$ as in (5.10) and (5.12).
@@
Theorem 5.3 immediately recovers the following property from [29, Thm. 4.3].
@@
So far we have avoided defining a skew-symmetric analogue of the ideal $I_{w}$ from (4.3). Such an ideal $I^{\mathrm{ss}}_{w}\subseteq\mathbb{K}[\mathsf{Mat}^{\mathrm{ss}}_{n\times n}]$ is introduced for each $w\in I_{{\textsf{FPF}}}^{(n)}$ in [29, Def. 3.13]. We omit an explicit description here, except to state that this ideal $I^{\mathrm{ss}}_{w}$ can be generated a set of Pfaffians of certain submatrices of the skew-symmetric matrix of variables $\mathcal{U}^{\mathrm{ss}}=\begin{bmatrix}0&-u_{ij}\\ u_{ij}&0\end{bmatrix}$.
The simplest Pfaffian generating set for $I^{\mathrm{ss}}_{w}$ given in [29, Def. 3.13] is not a Gröbner basis [29, Ex. 3.28]. However, there is a more complicated set of Pfaffians $G_{w}^{\mathrm{ss}}\subseteq I^{\mathrm{ss}}_{w}$ described in [29, Thm. 4.6] for which it turns out that $J^{\mathrm{ss}}_{w}\subseteq\langle\operatorname{init}(g):g\in G_{w}^{\mathrm{ss}}\rangle$ [29, Lems. 3.35 and 3.36].
It can be shown that $\mathsf{MSV}^{\mathrm{ss}}_{w}$ is the zero locus of $I^{\mathrm{ss}}_{w}$ and so $G_{w}^{\mathrm{ss}}\subseteq I^{\mathrm{ss}}_{w}\subseteq I(\mathsf{MSV}^{\mathrm{ss}}_{w})$ [29, Thm. 3.17]. Therefore, Corollary 3.6 implies that $G_{w}^{\mathrm{ss}}$ is a Gröbner basis for $I(\mathsf{MSV}^{\mathrm{ss}}_{w})$, so $I^{\mathrm{ss}}_{w}=\langle G_{w}^{\mathrm{ss}}\rangle=I(\mathsf{MSV}^{\mathrm{ss}}_{w})$ is prime (as $\mathsf{MSV}^{\mathrm{ss}}_{w}$ is irreducible) and $J^{\mathrm{ss}}_{w}=\operatorname{init}(I^{\mathrm{ss}}_{w})$. These facts recover [29, Thms. 4.5 and 4.6].
@@ 6 Symmetric matrix Schubert varieties
$\mathsf{MSC}^{\mathrm{sym}}_{w}$ =$=\left\{M\in\mathsf{Mat}^{\mathrm{sym}}_{n\times n}:\text{$\operatorname{rank}M_{[i][j]}=\operatorname{rank}w_{[i][j]}$ for all $(i,j)\in[n]\times[n]$}\right\},$ (6.1)
$\mathsf{MSV}^{\mathrm{sym}}_{w}$ =$=\left\{M\in\mathsf{Mat}^{\mathrm{sym}}_{n\times n}:\text{$\operatorname{rank}M_{[i][j]}\leq\operatorname{rank}w_{[i][j]}$ for all $(i,j)\in[n]\times[n]$}\right\},$
@@
As with $\mathsf{MSV}^{\mathrm{ss}}_{w}$, the varieties $\mathsf{MSV}^{\mathrm{sym}}_{w}$ include many familiar classes of symmetric matrices, whose ideals have been previously studied (see, e.g., [15, 25]). Questions about the Gröbner geometry of symmetric matrix Schubert varieties are less well-understood compared to our two previous cases. We know of only a few relevant references [6, 7, 10, 16], which supply answers in special cases.
@@ 6.1 An incomplete transition system
Fix an element $w\in\mathsf{Mat}^{\mathrm{sym}}_{n\times n}$. At least when $\operatorname{\mathsf{char}}(\mathbb{K})\neq 2$, the variety $\mathsf{MSV}^{\mathrm{sym}}_{w}$ is irreducible and equal to the Zariski closure of $\mathsf{MSC}^{\mathrm{sym}}_{w}$ [1, Lem. 5.2]. The reference [1] works over $\mathbb{C}$, but the relevant arguments hold over any algebraically closed field with $\operatorname{\mathsf{char}}(\mathbb{K})\neq 2$.
@@
As observed in Example 3.9, the transition systems realizing Conjecture 6.3 will have to include non-radical ideals of $\mathbb{K}[\mathsf{Mat}^{\mathrm{sym}}_{n\times n}]$. We expect that these systems will still be finite, however, as the subset (6.6) is itself finite. Specifically, we can restrict $w$ in (6.6) to range over all symmetric $n\times n$ partial permutation matrices, since these matrices give rise to all possible rank tables for elements of $\mathsf{Mat}^{\mathrm{sym}}_{n\times n}$. This claim can be deduced (when $\mathbb{K}$ is an arbitrary field) from [26, Lem. 3.20], which shows that the rank table of any matrix coincides with the rank table of some partial permutation matrix, and the latter must be symmetric for its rank table to be symmetric.
@@
$I_{\infty}:=\{w\in S_{\infty}:w=w^{-1}\}\quad\text{and}\quad I_{\infty}^{(n)}:=\{w\in I_{\infty}:\mathrm{Des}_{R}(w)\subseteq[n]\}=I_{\infty}\cap S_{\infty}^{(n,n)}.$ (6.7)
Passing to the $n\times n$ partial permutation matrix gives a bijection from $I_{\infty}^{(n)}$ to symmetric $n\times n$ partial permutation matrices. This lets us define $\mathsf{MSV}^{\mathrm{sym}}_{w}$, $I^{\mathrm{sym}}_{w}$, and $J^{\mathrm{sym}}_{w}$ for $w\in I_{\infty}^{(n)}$. Then
$\left\{(I(\mathsf{MSV}^{\mathrm{sym}}_{w}),J^{\mathrm{sym}}_{w}):w\in\mathsf{Mat}^{\mathrm{sym}}_{n\times n}\right\}=\left\{(I(\mathsf{MSV}^{\mathrm{sym}}_{w}),J^{\mathrm{sym}}_{w}):w\in I_{\infty}^{(n)}\right\},$ (6.8)
and the second set is uniquely indexed since if $v,w\in I_{\infty}^{(n)}$ then $I(\mathsf{MSV}^{\mathrm{sym}}_{v})\subseteq I(\mathsf{MSV}^{\mathrm{sym}}_{w})$ if and only if $v\leq w$ in the Bruhat order of $S_{\infty}$ by [1, Lem. 3.5].
@@
The ideal $I(\mathsf{MSV}^{\mathrm{sym}}_{w})$ is maximal if and only if $\mathsf{MSV}^{\mathrm{sym}}_{w}$ is a point, which occurs only when $\mathsf{MSV}^{\mathrm{sym}}_{w}=\{0\}$. The index $w\in I_{\infty}^{(n)}$ for this case is $0_{n\times n}=(1,n+1)(2,n+2)\cdots(n,2n)$, and we have $I(\mathsf{MSV}^{\mathrm{sym}}_{0_{n\times n}})=I^{\mathrm{sym}}_{0_{n\times n}}=J^{\mathrm{sym}}_{0_{n\times n}}=\operatorname{init}(I(\mathsf{MSV}^{\mathrm{sym}}_{0_{n\times n}}))=\langle u_{ij}:i,j\in[n],\ i\geq j\rangle.$
We mention some interesting consequences of Conjecture 6.3. First, via Theorem 3.3, the conjecture would immediately imply that $J^{\mathrm{sym}}_{w}$ is the initial ideal of $I(\mathsf{MSV}^{\mathrm{sym}}_{w})$. Then it would follow by Corollary 3.6 that the minors listed in (6.3) form a Gröbner basis for $I(\mathsf{MSV}^{\mathrm{sym}}_{w})$. As these minors already generate $I^{\mathrm{sym}}_{w}$, this would prove the following statements:
@@
With some computer assistance, we can directly construct transition systems verifying Conjecture 6.3 for $n\leq 4$. The relevant calculations are explained in Section 3. Using the computer algebra system Macaulay2, we have also been able to check that Conjecture 6.6 holds when $n\leq 7$ for any choice of field $\mathbb{K}$. We can also computationally verify Conjecture 6.7 when $n\leq 5$ and $\mathbb{K}$ is any of the (not algebraically closed) fields $\mathbb{Q}$, $\mathbb{F}_{2}$, or $\mathbb{F}_{3}$.
Remark 6.8.
One application of the Gröbner basis computations for matrix Schubert varieties and their skew-symmetric counterparts is to determine primary decompositions of the initial ideals $J_{w}$ and $J^{\mathrm{ss}}_{w}$ for $I_{w}=I(\mathsf{MSV}_{w})$ and $I^{\mathrm{ss}}_{w}=I(\mathsf{MSV}^{\mathrm{ss}}_{w})$. These decompositions are intersections of monomial ideals indexed by certain pipe dreams of $w$; see [23, Thm. B] and [26, Thm. 4.15].
@@
If $y\in I_{\infty}^{(n)}$ then $J^{\mathrm{sym}}_{y}=\bigcap_{D}J^{\mathrm{sym}}_{D}$ where $J^{\mathrm{sym}}_{D}:=\left\langle u_{ij}^{m_{D}(i,j)}:(i,j)\in D\right\rangle$ and where $D$ runs over all involution pipe dreams for all $z\in I_{\infty}^{(n)}$ with $z\geq y$ in Bruhat order.
While this conjecture is simple to state, it is probably not in its optimal form, as the given decomposition is very redundant. It suffices to let $D$ run over the minimal elements of the poset on involution pipe dreams for $z\geq y$ defined by $D\leq E$ if $J^{\mathrm{sym}}_{D}\subseteq J^{\mathrm{sym}}_{E}$. If the exponents $m_{D}(i,j)$ in the definition of $J^{\mathrm{sym}}_{D}$ were replaced by $1$, these minimal elements would be just the involution pipe dreams for $y$—this does seem to hold if $y$ is non-crossing (i.e., $3412$-avoiding), but not in general.

### v3

@@ 5 Skew-symmetric matrix Schubert varieties
$\mathsf{MSC}^{\mathrm{ss}}_{w}$ $=\left\{M\in\mathsf{Mat}^{\mathrm{ss}}_{n\times n}:\text{$\operatorname{rank}M_{[i][j]}=\operatorname{rank}w_{[i][j]}$ for all $(i,j)\in[n]\times[n]$}\right\},$ (5.1)
$\mathsf{MSV}^{\mathrm{ss}}_{w}$ $=\left\{M\in\mathsf{Mat}^{\mathrm{ss}}_{n\times n}:\text{$\operatorname{rank}M_{[i][j]}\leq\operatorname{rank}w_{[i][j]}$ for all $(i,j)\in[n]\times[n]$}\right\},$
@@ 5.1 Another reverse lexicographic transition system
Unlike the story for ordinary matrix Schubert varieties, both of these containments can be strict [31, Ex. 3.12]. Therefore, to get a skew-symmetric version of Theorem 4.3, we cannot use the right hand side of (5.4) as a skew-symmetric analogue of $J_{w}$.
Instead, we consider a more elaborate construction from [31]. Suppose $A=\{a_{0}<a_{1}<\dots<a_{r}\}$ and $B=\{b_{0}<b_{1}<\dots<b_{r}\}$ are two sets of $r+1$ positive integers. Define
@@
Then let $A\boxplus B:=\{(i,j)\in A\odot B:i\geq j\}\cup\{(i,j)\in B\odot A:i\geq j\}$ and set
@@
This can be shown in a self-contained way, but the elementary argument is fairly tedious. Alternatively, a special case of [31, Lem 3.25] asserts that $u^{\mathrm{ss}}_{AB}$ divides $u_{ij}$ if and only if $\operatorname{adiag}(\mathcal{X}_{AB})$ divides $x_{ij}x_{ji}$, which occurs either when $A=B=\{i,j\}$ or $\{A,B\}=\{\{i\},\{j\}\}$. ∎
@@ 5.2 Finite indexing set
For each $w\in\mathcal{I}_{\textsf{FPF}}$ define $\mathrm{ss}_{n}(w)\in\mathsf{Mat}^{\mathrm{ss}}_{n\times n}$ to be the matrix whose entry in position $(i,j)$ is $1$ if $i<j=w(i)$, 1-1 if $i>j=w(i)$, or else zero. Then each $\mathsf{B}_{n}$-orbit in $\mathsf{Mat}^{\mathrm{ss}}_{n\times n}$ contains an element of the form $\mathrm{ss}_{n}(w)$ for a unique $w\in\mathcal{I}_{\textsf{FPF}}^{(n)}$ by results in [4]; it is not hard to show that this holds even when $\mathbb{K}$ is not algebraically closed. Given $w\in\mathcal{I}_{\textsf{FPF}}^{(n)}$ we are therefore motivated to define
@@
$\mathcal{T}^{\mathrm{ss}}=\left\{(I(\mathsf{MSV}^{\mathrm{ss}}_{w}),J^{\mathrm{ss}}_{w}):w\in\mathsf{Mat}^{\mathrm{ss}}_{n\times n}\right\}=\left\{(I(\mathsf{MSV}^{\mathrm{ss}}_{w}),J^{\mathrm{ss}}_{w}):w\in\mathcal{I}_{\textsf{FPF}}^{(n)}\right\}$
and the second set is uniquely indexed, since if $v,w\in\mathcal{I}_{\textsf{FPF}}^{(n)}$ then $I(\mathsf{MSV}^{\mathrm{ss}}_{v})\subseteq I(\mathsf{MSV}^{\mathrm{ss}}_{w})$ if and only if $v\leq w$ in a certain Bruhat order on $\mathcal{I}_{\textsf{FPF}}$ [31, Prop. 3.11].
@@
As in the classical case, the ideal $I(\mathsf{MSV}^{\mathrm{ss}}_{w})$ is maximal if and only if $\mathsf{MSV}^{\mathrm{ss}}_{w}$ is a point, which occurs only when $\mathsf{MSV}^{\mathrm{ss}}_{w}=\{0\}$. The index $w\in\mathcal{I}_{\textsf{FPF}}^{(n)}$ for this case is
$0^{\mathrm{ss}}_{n\times n}:=(1,n+1)(2,n+2)\cdots(n,2n)(2n+1,2n+2)(2n+3,2n+4)\cdots\in\mathcal{I}_{\textsf{FPF}}^{(n)}$
@@
Suppose $w\in\mathcal{I}_{\textsf{FPF}}^{(n)}$ and $(i,j)\in[n]\times[n]$ has $i>j$. Then $(i,j)\in\mathsf{dom}(w)$ if and only if $u_{ij}\in I(\mathsf{MSV}^{\mathrm{ss}}_{w})$ if and only if $u_{ij}\in J^{\mathrm{ss}}_{w}$.
@@
Following [17, §4.1], we define the fpf-involution length of $w\in\mathcal{I}_{\textsf{FPF}}$ to be
@@
Write $u\lessdot_{\textsf{FPF}}v$ for $u,v\in\mathcal{I}_{\textsf{FPF}}$ if $v=t_{ij}\cdot u\cdot t_{ij}$ for any positive integers $i<j$ with $\ell_{\textsf{FPF}}(v)=\ell_{\textsf{FPF}}(u)+1$. For a description of $\lessdot_{\textsf{FPF}}$ in terms of the cycles of $u$ and $v$, see [17, Prop. 4.9].
Suppose $0^{\mathrm{ss}}_{n\times n}\neq w\in\mathcal{I}_{\textsf{FPF}}^{(n)}$ so that $I(\mathsf{MSV}^{\mathrm{ss}}_{w})$ is not a maximal ideal and $\mathsf{MSV}^{\mathrm{ss}}_{w}\neq\{0\}$. Then $\mathsf{dom}(w)$, which is invariant under transpose, cannot contain every position in $n]\times[n]$, so there must exist an outer corner $(p,q)$ of $\mathsf{dom}(w)$ with $n\geq p\geq q\geq 1$. As explained in [31, §4.1], all outer corners $(p,q)$ of $\mathsf{dom}(w)$ have $w(p)=q$, so as $w$ has no fixed points we can assume that $p>q$.
@@
It is known that $\varnothing\subsetneq\mathcal{C}^{\mathrm{ss}}(w)\subseteq\mathcal{I}_{\textsf{FPF}}^{(n)}$ by [31, Lem. 4.8]. Finally, let
@@
The claim that $J^{\mathrm{ss}}_{w}\subseteq\operatorname{init}(I(\mathsf{MSV}^{\mathrm{ss}}_{w}))$ holds by [31, Lem 3.37 and Thm. 3.17]. The other parts of axioms (T1) and (T2) in Definition 3.1 were checked in Sections 5.2 and 5.3.
Fix $w\in\mathcal{I}_{\textsf{FPF}}^{(n)}$ with $I(\mathsf{MSV}^{\mathrm{ss}}_{w})$ not maximal, and then choose an outer corner $(p,q)$ of $\mathsf{dom}(w)$ with $n\geq p>q\geq 1$. The identities required for axiom (T3) when $(I,J)=(I(\mathsf{MSV}^{\mathrm{ss}}_{w}),J^{\mathrm{ss}}_{w})$ are supplied in [31, Lem. 4.2] if we define $f=u_{pq}\notin I(\mathsf{MSV}^{\mathrm{ss}}_{w})$ and $\Phi$ as in (5.11) and (5.13).
@@
Theorem 5.3 immediately recovers the following property from [31, Thm. 4.3].
@@
So far we have avoided defining a skew-symmetric analogue of the ideal $I_{w}$ from (4.3). Such an ideal $I^{\mathrm{ss}}_{w}\subseteq\mathbb{K}[\mathsf{Mat}^{\mathrm{ss}}_{n\times n}]$ is introduced for each $w\in\mathcal{I}_{\textsf{FPF}}^{(n)}$ in [31, Def. 3.13]. We omit an explicit description here, except to state that this ideal $I^{\mathrm{ss}}_{w}$ can be generated a set of Pfaffians of certain submatrices of the skew-symmetric matrix of variables $\mathcal{U}^{\mathrm{ss}}=\begin{bmatrix}0&-u_{ij}\\ u_{ij}&0\end{bmatrix}$.
The simplest Pfaffian generating set for $I^{\mathrm{ss}}_{w}$ given in [31, Def. 3.13] is not a Gröbner basis [31, Ex. 3.28]. However, there is a more complicated set of Pfaffians $G_{w}^{\mathrm{ss}}\subseteq I^{\mathrm{ss}}_{w}$ described in [31, Thm. 4.6] for which it turns out that $J^{\mathrm{ss}}_{w}\subseteq\langle\operatorname{init}(g):g\in G_{w}^{\mathrm{ss}}\rangle$ [31, Lems. 3.35 and 3.36].
It can be shown that $\mathsf{MSV}^{\mathrm{ss}}_{w}$ is the zero locus of $I^{\mathrm{ss}}_{w}$ and so $G_{w}^{\mathrm{ss}}\subseteq I^{\mathrm{ss}}_{w}\subseteq I(\mathsf{MSV}^{\mathrm{ss}}_{w})$ [31, Thm. 3.17]. Therefore, Corollary 3.6 implies that $G_{w}^{\mathrm{ss}}$ is a Gröbner basis for $I(\mathsf{MSV}^{\mathrm{ss}}_{w})$, so $I^{\mathrm{ss}}_{w}=\langle G_{w}^{\mathrm{ss}}\rangle=I(\mathsf{MSV}^{\mathrm{ss}}_{w})$ is prime (as $\mathsf{MSV}^{\mathrm{ss}}_{w}$ is irreducible) and $J^{\mathrm{ss}}_{w}=\operatorname{init}(I^{\mathrm{ss}}_{w})$. These facts recover [31, Thms. 4.5 and 4.6].
@@ 6 Symmetric matrix Schubert varieties
$\mathsf{MSC}^{\mathrm{sym}}_{w}$ $=\left\{M\in\mathsf{Mat}^{\mathrm{sym}}_{n\times n}:\text{$\operatorname{rank}M_{[i][j]}=\operatorname{rank}w_{[i][j]}$ for all $(i,j)\in[n]\times[n]$}\right\},$ (6.1)
$\mathsf{MSV}^{\mathrm{sym}}_{w}$ $=\left\{M\in\mathsf{Mat}^{\mathrm{sym}}_{n\times n}:\text{$\operatorname{rank}M_{[i][j]}\leq\operatorname{rank}w_{[i][j]}$ for all $(i,j)\in[n]\times[n]$}\right\},$
@@
As with $\mathsf{MSV}^{\mathrm{ss}}_{w}$, the varieties $\mathsf{MSV}^{\mathrm{sym}}_{w}$ include many familiar classes of symmetric matrices, whose ideals have been previously studied (see, e.g., [15, 27]). Questions about the Gröbner geometry of symmetric matrix Schubert varieties are less well-understood. We know of only a few relevant references [6, 7, 10, 16], which supply answers in special cases.
@@ 6.1 An incomplete transition system
[…]
@@
As observed in Example 3.9, the transition systems realizing Conjecture 6.3 will have to include non-radical ideals of $\mathbb{K}[\mathsf{Mat}^{\mathrm{sym}}_{n\times n}]$. We expect that these systems will still be finite, however.
One can see that the set $\{(I(\mathsf{MSV}^{\mathrm{sym}}_{w}),J^{\mathrm{sym}}_{w}):w\in\mathsf{Mat}^{\mathrm{sym}}_{n\times n}\}$ is itself finite by noting that the set is unchanged when we restrict $w$ to range over all symmetric $n\times n$ partial permutation matrices, since these matrices give rise to all possible rank tables for elements of $\mathsf{Mat}^{\mathrm{sym}}_{n\times n}$. The latter claim can be deduced (when $\mathbb{K}$ is an arbitrary field) from [28, Lem. 3.20], which shows that the rank table of any matrix coincides with the rank table of some partial permutation matrix, and the latter must be symmetric for its rank table to be symmetric.
@@
$\mathcal{I}_{\infty}:=\{w\in S_{\infty}:w=w^{-1}\}\quad\text{and}\quad\mathcal{I}_{\infty}^{(n)}:=\{w\in\mathcal{I}_{\infty}:\mathrm{Des}_{R}(w)\subseteq[n]\}=\mathcal{I}_{\infty}\cap S_{\infty}^{(n,n)}.$ (6.7)
Passing to the $n\times n$ partial permutation matrix gives a bijection from $\mathcal{I}_{\infty}^{(n)}$ to symmetric $n\times n$ partial permutation matrices. This lets us define $\mathsf{MSV}^{\mathrm{sym}}_{w}$, $I^{\mathrm{sym}}_{w}$, and $J^{\mathrm{sym}}_{w}$ for $w\in\mathcal{I}_{\infty}^{(n)}$. Then
$\left\{(I(\mathsf{MSV}^{\mathrm{sym}}_{w}),J^{\mathrm{sym}}_{w}):w\in\mathsf{Mat}^{\mathrm{sym}}_{n\times n}\right\}=\left\{(I(\mathsf{MSV}^{\mathrm{sym}}_{w}),J^{\mathrm{sym}}_{w}):w\in\mathcal{I}_{\infty}^{(n)}\right\},$ (6.8)
and the second set is uniquely indexed since if $v,w\in\mathcal{I}_{\infty}^{(n)}$ then $I(\mathsf{MSV}^{\mathrm{sym}}_{v})\subseteq I(\mathsf{MSV}^{\mathrm{sym}}_{w})$ if and only if $v\leq w$ in the Bruhat order of $S_{\infty}$ by [1, Lem. 3.5].
@@
The ideal $I(\mathsf{MSV}^{\mathrm{sym}}_{w})$ is maximal if and only if $\mathsf{MSV}^{\mathrm{sym}}_{w}$ is a point, which occurs only when $\mathsf{MSV}^{\mathrm{sym}}_{w}=\{0\}$. The index $w\in\mathcal{I}_{\infty}^{(n)}$ for this case is $0_{n\times n}=(1,n+1)(2,n+2)\cdots(n,2n)$, and we have $I(\mathsf{MSV}^{\mathrm{sym}}_{0_{n\times n}})=I^{\mathrm{sym}}_{0_{n\times n}}=J^{\mathrm{sym}}_{0_{n\times n}}=\operatorname{init}(I(\mathsf{MSV}^{\mathrm{sym}}_{0_{n\times n}}))=\langle u_{ij}:i,j\in[n],\ i\geq j\rangle.$
We mention some interesting consequences of Conjecture 6.3. First, via Theorem 3.3, the conjecture would immediately imply that $J^{\mathrm{sym}}_{w}$ is the initial ideal of $I(\mathsf{MSV}^{\mathrm{sym}}_{w})$. Then it would follow by Corollary 3.6 that the minors listed in (6.3) form a Gröbner basis for $I(\mathsf{MSV}^{\mathrm{sym}}_{w})$. As these minors already generate $I^{\mathrm{sym}}_{w}$, this would prove the following:
@@
One application of the Gröbner basis computations for matrix Schubert varieties and their skew-symmetric counterparts is to determine primary decompositions of the initial ideals $J_{w}$ and $J^{\mathrm{ss}}_{w}$ for $I_{w}=I(\mathsf{MSV}_{w})$ and $I^{\mathrm{ss}}_{w}=I(\mathsf{MSV}^{\mathrm{ss}}_{w})$. These decompositions are intersections of monomial ideals indexed by certain pipe dreams of $w$; see [25, Thm. B] and [28, Thm. 4.15].
[…]
Computations for $n\leq 10$ support the following conjecture. Recall that $\mathcal{I}_{n}\subsetneq\mathcal{I}_{\infty}^{(n)}$ is the set of involutions in $S_{n}$, corresponding to the symmetric $n\times n$ partial permutation matrices of full rank.
@@
If $w\in\mathcal{I}_{n}$ then $J^{\mathrm{sym}}_{w}=\bigcap_{D}J^{\mathrm{sym}}_{D}$ where $J^{\mathrm{sym}}_{D}:=\left\langle u_{ij}^{m_{D}(i,j)}:(i,j)\in D\right\rangle$ and where $D$ runs over all involution pipe dreams for all $z\in\mathcal{I}_{n}$ with $z\geq w$ in Bruhat order.
Replacing $\mathcal{I}_{n}$ by $\mathcal{I}_{\infty}^{(n)}$ in this conjecture would give the most natural symmetric analogue of [31, Thm. 4.15]. However, this broader claim is false, and to decompose $J^{\mathrm{sym}}_{w}$ when $w$ corresponds to a symmetric $n\times n$ partial permutation matrix that is not invertible, a more involved statement is required.
We also mention that when $w\in\mathcal{I}_{n}$, the decomposition given in Conjecture 6.8 is very redundant. It suffices to let $D$ run over the minimal elements of the poset on involution pipe dreams for $z\geq w$ defined by $D\leq E$ if $J^{\mathrm{sym}}_{D}\subseteq J^{\mathrm{sym}}_{E}$. If the exponents $m_{D}(i,j)$ in the definition of $J^{\mathrm{sym}}_{D}$ were replaced by $1$, then these minimal elements would be just the involution pipe dreams for $w$. This does seem to hold if $w$ is non-crossing in the sense of being $3412$-avoiding, but is not true in general.